# Cost-Driven Representation Learning for Linear Quadratic Gaussian Control: Part I

### v3

@@ 1 Introduction
Notation. We use 0 (resp. $1$) to denote either the scalar or a matrix consisting of all zeros (resp. all ones); we use $I$ to denote an identity matrix. The dimension, when emphasized, is specified in subscripts, e.g., $0_{d_{x}\times d_{x}},1_{d_{x}},I_{d_{x}}$. Let $\mathbb{I}_{S}$ denote the indicator function for set SS and $\mathbb{I}_{S}(A)$ apply to matrix $A$ elementwise. For some positive semidefinite $P$, we define $\|v\|_{P}:=(v^{\top}Pv)^{1/2}$. Semicolon “;” denotes stacking vectors or matrices vertically. For a collection of $d$-dimensional vectors $(v_{t})_{t=i}^{j}$, let $v_{i:j}:=[v_{i};v_{i+1};\ldots;v_{j}]\in\mathbb{R}^{d(j-i+1)}$ denote the concatenation along the column. For random variable $\eta$, let $\|\eta\|_{\psi_{\beta}}$ denote its $\beta$-sub-Weibull norm, a special case of Orlicz norms (Zhang and Wei, 2022), with $\beta=1,2$ corresponding to subexponential and sub-Gaussian norms. For matrix $A$, let $\sigma_{i}(A),\sigma_{\min}(A),\sigma_{\min}^{+}(A),\sigma_{\max}(A)$ denote its $i$th largest, minimum, minimum positive, maximum singular values, respectively. $\|A\|_{2},\|A\|_{F},\|A\|_{\ast}$ denote the operator (induced by vector $2$-norms), Frobenius, nuclear norms of matrix $A$, respectively. $\left\langle\cdot,\cdot\right\rangle_{F}$ denotes the Frobenius inner product between matrices. For square matrix $A$, let $\lambda_{\min}(A)$ be its minimum eigenvalue. The Kronecker, symmetric Kronecker, and Hadamard products between matrices are denoted by $\otimes$”, $\otimes_{s}$” and $\odot$”, respectively. $\mathrm{vec}(\cdot)$ and $\mathrm{svec}(\cdot)$ denote flattening a matrix and a symmetric matrix by stacking their columns; $\mathrm{svec}(\cdot)$ does not repeat the off-diagonal elements, but scales them by $\sqrt{2}$ (Schacke, 2004). Let $\textnormal{diag}(\cdot)$ denote the block diagonal matrix formed by the matrices inside the parentheses. For $\Sigma_{i=a}^{b}$, we define the sum to be zero if the lower index $a$ is greater than the upper index $b$.
@@ 2 Problem setup
Assumption 1 is standard in controlling LTV systems (Zhou and Zhao, 2017; Hu and Mitsui, 2017; Minasyan et al., 2021), satisfied by a stable LTI system. It essentially says that zero control is a stabilizing controller, and can be potentially relaxed to the assumption of being given a stabilizing controller as in Lale et al. (2020); Simchowitz et al. (2020), where one can excite the system using the stabilizing controller plus Gaussian random noises.
@@ 3 Methodology: Cost-driven state representation learning
Subramanian et al. (2020) propose to optimize a simple combination of cost and transition prediction errors to learn the approximate information state (Subramanian et al., 2020). That is, we parameterize a state representation function by matrices $(M_{t})_{t=0}^{T}$ and a latent model by matrices $((A_{t},B_{t})_{t=0}^{T-1},(Q_{t})_{t=0}^{T})$ and then solve
@@ 3.1 Learning the state representation function
6:Return: state representation estimate $(\hat{M}_{t})_{t=0}^{T}$ and latent state estimates $(\hat{z}_{t}^{(i)})_{t=0,i=1}^{T,n}$
@@ 4 Theoretical guarantees
Given an unknown LQG control problem defined by (2.1) and (2.2), under Assumptions 1234 and 5, for a given $p\in(0,1)$, if we run Algorithm 1 for $n\geq\mathrm{poly}(T,d_{x},d_{y},d_{u},\log(1/p))$, then with probability at least 1p1-p, state representation function $(\hat{M}_{t})_{t=0}^{T}$ is $\mathrm{poly}(\ell,d_{x},d_{y},d_{u},\log(\ell/p))n^{-1/4}$-optimal in the first $\ell$ steps, and $\mathrm{poly}(\nu^{-1},T,d_{x},d_{y},d_{u},\log(T/p))n^{-1/2}$-optimal in the next $(T-\ell)$ steps. In addition, the overall output policy $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T-1}$ satisties
@@
Theorem 1 provides finite-sample guarantees for both the state representation function $(\hat{M}_{t})_{t=0}^{T}$ and the overall policy $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T}$. From Theorem 1, we observe a separation of the sample complexities before and after time step $\ell$ for the state representation function, resulting from the loss of the full-rankness of $(\mathrm{Cov}(z^{\ast}_{t}))_{t=0}^{\ell-1}$ and $(M^{\ast}_{t})_{t=0}^{\ell-1}$. Specifically, quadratic regression guarantees that $\hat{N}_{t}$ converges to $N^{\ast}_{t}$ at a rate of $n^{-1/2}$ for all $0\leq t\leq T$ (Lemma 2). Before time step $\ell$, $\hat{M}_{t}$ suffers a square root decay of the rate $n^{-1/4}$ because $M^{\ast}_{t}$ may not have rank $d_{x}$ (Lemma 4). Since $(\hat{z}_{t})_{t=0}^{\ell-1}$ may not have full-rank covariances, $(A^{\ast}_{t})_{t=0}^{\ell-1}$ are only recovered partially. As a result, $(\hat{K}_{t})_{t=0}^{\ell-1}$ may not stabilize $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$. This issue poses a significant challenge for the analysis and significantly worsens dependence on $\ell$ in the policy suboptimality gap (Lemma 7), which implies that if $n$ is not large enough, the policy may be inferior to zero control in the first $\ell$ steps, as system $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$ is uniformly exponential stable (Assumption 1), while zero control has a suboptimality gap linear in $\ell$. After time step $\ell$, $\hat{M}_{t}$ retains the $n^{-1/2}$ sample complexity, from which the same order of sample complexity guarantee for $(\hat{A}_{t},\hat{B}_{t})$ follows, resulting in the $\mathrm{poly}(T)n^{-1}$ term in the policy suboptimality gap.
@@ 4.2 Quadratic regression bound
As noted in §3.1, the quadratic regression can be converted to linear regression using $\|h\|_{P}^{2}=\left\langle hh^{\top},P\right\rangle_{F}=\left\langle\mathrm{svec}(hh^{\top}),\mathrm{svec}(P)\right\rangle$. To analyze this linear regression with an intercept, we need the following lemma. We note that a similar lemma without considering the intercept has been proved in Jadbabaie et al. (2021, Proposition 1).
@@
Then, with the similar concentration arguments to those in Jadbabaie et al. (2021, Appendix A.1), we can show that with probability at least 1p1-p,
@@
$\|F^{\ast}-F\|_{F}^{2}=\sum\nolimits_{i=1}^{n}\|(h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top}-h^{(i)}(h^{(i)})^{\top}\|_{F}^{2}=\mathcal{O}(\epsilon^{2}d\|\Sigma_{\ast}\|_{2}n\log^{2}(n/p)).$
It follows that
@@
Using arguments similar to those in Mhammedi et al. (2020, Section B.2.13), we have
@@
Since $\xi$ is a vector of zero-mean subexponential variables with subexponential norms bounded by $E$, $\|\xi\|=\mathcal{O}(En^{1/2}\log(n/p))$. Hence,
@@
To bound $\|(\overline{F}^{\ast})^{\top}\xi\|=\|\textnormal{diag}(\Phi^{\ast},1)\overline{F}_{0}^{\top}\xi\|$, note that $\|\overline{F}_{0}^{\top}\xi\|=\|\sum_{i=1}^{n}\overline{f}_{0}^{(i)}e^{(i)}\|$. Consider the $j$th component in the summation $\sum_{i=1}^{n}[\overline{f}_{0}^{(i)}]_{j}(e^{\prime})^{(i)}$. Recall that $f_{0}=\mathrm{svec}(h_{0}h_{0}^{\top})$, so $\overline{f}_{0}]_{j}$ is either the square of a standard Gaussian random variable, $\sqrt{2}$ times the product of two independent standard Gaussian random variables, or one. Hence, $f_{0}]_{j}$ is subexponential with mean and $\|[f_{0}]_{j}\|_{\psi_{1}}$ both bounded by $\mathcal{O}(1)$. As a result, the product $f_{0}]_{j}e$ is $\frac{1}{2}$-sub-Weibull, with the sub-Weibull norm being $\mathcal{O}(E)$. By Hao et al. (2019, Theorem 3.1),
@@ 4.3 Matrix factorization bound
Given two $m\times n$ matrices $A,B$, we are interested in bounding $\min_{S^{\top}S=I}\|SA-B\|_{F}$ using $\|A^{\top}A-B^{\top}B\|_{F}$. The minimum problem is known as the orthogonal Procrustes problem, solved in Schoenemann (1964); Schönemann (1966). Specifically, the minimum is attained at $S=UV^{\top}$, where $U\Sigma V^{\top}=BA^{\top}$ is its singular value decomposition.
If $m\leq n$ and $\mathrm{rank}(A)=m$, then the following lemma from Tu et al. (2016) establishes that the distance between $A$ and $B$ is of the same order of $\|A^{\top}A-B^{\top}B\|_{F}$.
Lemma 3 (Tu et al. (2016, Lemma 5.4)).
@@ 4.4 Perturbed linear regression bound
By Wainwright (2019, Theorem 6.1), the Gaussian ensemble $G$ satisfies that with probability at least 1p1-p,
@@
where we consider $\epsilon_{x}$ and $\epsilon_{y}$ as quantities much smaller than one such that terms like $\epsilon_{x}^{2},\epsilon_{x}\epsilon_{y}$ are absorbed into $\epsilon_{x},\epsilon_{y}$. It remains to control $\|G^{\dagger}E\|_{2}$. Mhammedi et al. (2020, Section B.2.11) prove via a covering number argument that with probability at least 1p1-p,
@@ 4.4.1 Proof of Claim 1
Similar to the proof of Lemma 5, by Mhammedi et al. (2020, Section B.2.11), with probability at least 1p1-p,
@@ 4.5.3 Proof of Claim 4
The following lemma is the counterpart of Fazel et al. (2018, Lemma 10) in the LTV setting with system noises.
@@
Mania et al. (2019) have studied this problem in the infinite-horizon LTI setting; here we extend their result to the finite-horizon LTV setting.
We adopt the following compact formulation of a finite-horizon LTV system, as introduced in Zhang et al. (2021, Section 3), to reduce our setting to the infinite-horizon LTI one:
@@
Let $\hat{P}=\textnormal{diag}(\hat{P}_{0},\ldots,\hat{P}_{T})$ be the optimal cumulative cost matrices in system $(\hat{A},\hat{B},\hat{Q},R^{\ast})$ by the RDE (2.6). With a slight abuse of notation, define $K^{\ast}:=[\textnormal{diag}(K^{\ast}_{0},\ldots,K^{\ast}_{T-1}),0_{d_{u}T\times d_{x}}]$, where $(K^{\ast}_{t})_{t=0}^{T-1}$ is the optimal controller in system $(A^{\ast},B^{\ast},Q^{\ast},R^{\ast})$, and define $\hat{K}$ similarly for system $(\hat{A},\hat{B},\hat{Q},R^{\ast})$. By definition, $A^{\ast}$ is stable (hence, stabilizable), $(A^{\ast},Q^{\ast})$ is observable in the sense of LTI systems, and $R^{\ast}\succ 0$. Therefore, by Mania et al. (2019, Propositions 1 and 2), there exists a dimension-free constant $\varepsilon_{0}>0$ with $\varepsilon_{0}^{-1}$ depending polynomially on problem parameters such that as long as $\varepsilon\leq\varepsilon_{0}$,
@@
Finally, we are ready to present the main lemma of this section below, which combines Lemmas 7 to 9 to provide end-to-end guarantees for the certainty equivalent optimal controller in linear quadratic control with both rank-deficient and full-rank stages.
@@
For any feedback gains $(K_{t})_{t=0}^{T-1}$, let $J(K)$ denote the expected cumulative cost in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1}$, $(Q^{\ast}_{t})_{t=0}^{T})$ under state feedback control $u_{t}=K_{t}x_{t}$ for $t\geq 0$, and $J^{\delta}(K)$ denote the expected cumulative cost in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ under control $u_{t}=K_{t}(x_{t}+\delta_{t})$, where $\delta_{t}$ is a perturbation, for $t\geq 0$. Under control $u_{t}=K_{t}(x_{t}+\delta_{t})$ for all $t\geq 0$, let $\Xi_{t}(K):=\mathrm{Cov}(x_{t})$ in the system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$, and assume that $\delta_{t}$ follows an arbitrary zero-mean Gaussian distribution, can be correlated with $x_{t}$, and satisfies $\|\mathrm{Cov}(\delta_{t})\|_{2}^{1/2}\leq\varphi_{1}\max_{\tau\leq t}\|\Xi_{\tau}(K)\|_{2}^{1/2}$ for $0\leq t\leq\ell-1$ and $\|\mathrm{Cov}(\delta_{t})\|_{2}^{1/2}\leq\varphi_{2}\max_{\tau\leq t}\|\Xi_{\tau}(K)\|_{2}^{1/2}$ for $t\geq\ell$. There exist dimension-free constants $\kappa,a_{2},a_{3},a_{4}>1$ that depend polynomially on $c$ and other problem parameters and dimension-free constant $\varepsilon_{0}>0$ with $\varepsilon_{0}^{-1}$ depending polynomially on problem parameters, such that under the conditions that 1) $\varepsilon_{1},\varphi_{1}$ are small enough to ensure $\|\hat{B}_{t}\|_{2},\|\hat{Q}_{t}\|_{2}$ are of order $\mathcal{O}(1)$, $\lambda_{\min}(\hat{Q}_{t})=\Omega(1)$ for $0\leq t\leq\ell-1$ and $\varepsilon_{1}+\varphi_{1}=\mathcal{O}((\kappa a_{2}+a_{3}^{2})^{-T})$, and that 2) $\varepsilon_{2},\varphi_{2}$ are small enough to ensure $\varepsilon_{2}\leq\varepsilon_{0}$ and $\varepsilon_{2}+\varphi_{2}=\mathcal{O}(\varepsilon_{1}+\varphi_{1})=\mathcal{O}({T}^{-1})$, we have
@@
$J_{2}(\hat{K})=\mathbb{E}\big{[}\|x_{\ell}^{\delta}\|_{P_{\ell}^{\hat{K}}}^{2}\big{]}+\sum\nolimits_{t=\ell}^{T-1}\mathbb{E}[w_{t}^{\top}P_{t+1}^{\hat{K}}w_{t}],$

### v4

@@ 1 Introduction
Notation. We use 0 (resp. $1$) to denote either the scalar or a matrix consisting of all zeros (resp. all ones); we use $I$ to denote an identity matrix. The dimension, when emphasized, is specified in subscripts, e.g., $0_{d_{x}\times d_{x}},1_{d_{x}},I_{d_{x}}$. Let $\mathbb{I}_{S}$ denote the indicator function for set $S$ and $\mathbb{I}_{S}(A)$ apply to matrix $A$ elementwise. For some positive semidefinite $P$, we define $\|v\|_{P}:=(v^{\top}Pv)^{1/2}$. Semicolon “;” denotes stacking vectors or matrices vertically. For a collection of $d$-dimensional vectors $(v_{t})_{t=i}^{j}$, let $v_{i:j}:=[v_{i};v_{i+1};\ldots;v_{j}]\in\mathbb{R}^{d(j-i+1)}$ denote the concatenation along the column. For random variable $\eta$, let $\|\eta\|_{\psi_{\beta}}$ denote its $\beta$-sub-Weibull norm, a special case of Orlicz norms (Zhang and Wei, 2022), with $\beta=1,2$ corresponding to subexponential and sub-Gaussian norms. For matrix $A$, let $\sigma_{i}(A),\sigma_{\min}(A),\sigma_{\min}^{+}(A),\sigma_{\max}(A)$ denote its $i$th largest, minimum, minimum positive, maximum singular values, respectively. $\|A\|_{2},\|A\|_{F},\|A\|_{\ast}$ denote the operator (induced by vector $2$-norms), Frobenius, nuclear norms of matrix $A$, respectively. $\left\langle\cdot,\cdot\right\rangle_{F}$ denotes the Frobenius inner product between matrices. For square matrix $A$, let $\lambda_{\min}(A)$ be its minimum eigenvalue. The Kronecker, symmetric Kronecker, and Hadamard products between matrices are denoted by $\otimes$”, $\otimes_{s}$” and $\odot$”, respectively. $\mathrm{vec}(\cdot)$ and $\mathrm{svec}(\cdot)$ denote flattening a matrix and a symmetric matrix by stacking their columns; $\mathrm{svec}(\cdot)$ does not repeat the off-diagonal elements, but scales them by $\sqrt{2}$ (Schacke, 2004). Let $\textnormal{diag}(\cdot)$ denote the block diagonal matrix formed by the matrices inside the parentheses. For $\Sigma_{i=a}^{b}$, we define the sum to be zero if the lower index $a$ is greater than the upper index $b$.
@@ 2 Problem setup
Assumption 1 is standard in controlling LTV systems (Zhou and Zhao, 2017; Hu and Mitsui, 2017; Minasyan et al., 2021), satisfied by a stable LTI system. It essentially says that zero control is a stabilizing controller, and can be potentially relaxed to the assumption of being given a stabilizing controller as in (Lale et al., 2020; Simchowitz et al., 2020), where one can excite the system using the stabilizing controller plus Gaussian random noises.
@@ 3 Methodology: Cost-driven state representation learning
Subramanian et al. (2020) propose to optimize a simple combination of cost and transition prediction errors to learn what they call the approximate information state. That is, we parameterize a state representation function by matrices $(M_{t})_{t=0}^{T}$ and a latent model by matrices $((A_{t},B_{t})_{t=0}^{T-1},(Q_{t})_{t=0}^{T})$ and then solve
@@ 3.1 Learning the state representation function
6:Return: state representation function estimate $(\hat{M}_{t})_{t=0}^{T}$ and latent state estimates $(\hat{z}_{t}^{(i)})_{t=0,i=1}^{T,n}$
@@ 4 Theoretical guarantees and proofs
Given an unknown LQG control problem defined by (2.1) and (2.2), under Assumptions 1234 and 5, for a given $p\in(0,1)$, if we run CoReL (Algorithm 1) for $n\geq\mathrm{poly}(T,d_{x},d_{y},d_{u},\log(1/p))$, then with probability at least 1p1-p, state representation function $(\hat{M}_{t})_{t=0}^{T}$ is $\mathrm{poly}(\ell,d_{x},d_{y},d_{u},\log(\ell/p))n^{-1/4}$-optimal in the first $\ell$ steps, and $\mathrm{poly}(\nu^{-1},T,d_{x},d_{y},d_{u},\log(T/p))n^{-1/2}$-optimal in the next $(T-\ell)$ steps. In addition, the overall output policy $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T-1}$ satisties
@@
Theorem 1 provides finite-sample guarantees for both the state representation function $(\hat{M}_{t})_{t=0}^{T}$ and the overall policy $\hat{\pi}=(\hat{M}_{t},\hat{K}_{t})_{t=0}^{T}$. From Theorem 1, we observe a separation of the sample complexities before and after time step $\ell$ for the state representation function, resulting from the loss of the full-rankness of $(\mathrm{Cov}(z^{\ast}_{t}))_{t=0}^{\ell-1}$ and $(M^{\ast}_{t})_{t=0}^{\ell-1}$.
Specifically, quadratic regression guarantees that $\hat{N}_{t}$ converges to $N^{\ast}_{t}$ at a rate of $n^{-1/2}$ for all $0\leq t\leq T$ (Lemma 2). Before time step $\ell$, $\hat{M}_{t}$ suffers a square root decay of the rate $n^{-1/4}$ because $M^{\ast}_{t}$ may not have rank $d_{x}$ (Lemma 4). Since $(\hat{z}_{t})_{t=0}^{\ell-1}$ may not have full-rank covariances, $(A^{\ast}_{t})_{t=0}^{\ell-1}$ are only recovered partially. As a result, $(\hat{K}_{t})_{t=0}^{\ell-1}$ may not stabilize $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$. This issue poses a significant challenge for the analysis and significantly worsens dependence on $\ell$ in the policy suboptimality gap (Lemma 7), which implies that if $n$ is not large enough, the policy may be inferior to zero control in the first $\ell$ steps, as system $(A^{\ast}_{t},B^{\ast}_{t})_{t=0}^{\ell-1}$ is uniformly exponential stable (Assumption 1), while zero control has a suboptimality gap linear in $\ell$. After time step $\ell$, $\hat{M}_{t}$ retains the $n^{-1/2}$ sample complexity, from which the same order of sample complexity guarantee for $(\hat{A}_{t},\hat{B}_{t})$ follows, resulting in the $\mathrm{poly}(T)n^{-1}$ term in the policy suboptimality gap.
@@ 4.2 Quadratic regression bound
As noted in §3.1, the quadratic regression can be converted to linear regression using $\|h\|_{P}^{2}=\left\langle hh^{\top},P\right\rangle_{F}=\left\langle\mathrm{svec}(hh^{\top}),\mathrm{svec}(P)\right\rangle$. To analyze this linear regression with an intercept, we need the following lemma. We note that a similar lemma without considering the intercept has been proved in (Jadbabaie et al., 2021, Proposition 1).
@@
Then, with the similar concentration arguments to those in (Jadbabaie et al., 2021, Appendix A.1), we can show that with probability at least 1p1-p,
@@
$\|F^{\ast}-F\|_{F}^{2}=\sum\nolimits_{i=1}^{n}\|(h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top}-h^{(i)}(h^{(i)})^{\top}\|_{F}^{2}=\mathcal{O}(\epsilon^{2}d\|\Sigma_{\ast}\|_{2}n\log^{2}(n/p)),$
which implies that
@@
Using arguments similar to those in (Mhammedi et al., 2020, Section B.2.13), we have
@@
Since $\xi$ is a vector of zero-mean subexponential variables with subexponential norms bounded by $E$, $\|\xi\|=\mathcal{O}(En^{1/2}\log(n/p))$. Hence, we have
@@
To bound $\|(\overline{F}^{\ast})^{\top}\xi\|=\|\textnormal{diag}(\Phi^{\ast},1)\overline{F}_{0}^{\top}\xi\|$, note that $\|\overline{F}_{0}^{\top}\xi\|=\|\sum_{i=1}^{n}\overline{f}_{0}^{(i)}e^{(i)}\|$. Consider the $j$th component in the summation $\sum_{i=1}^{n}[\overline{f}_{0}^{(i)}]_{j}(e^{\prime})^{(i)}$. Recall that $f_{0}=\mathrm{svec}(h_{0}h_{0}^{\top})$, so $\overline{f}_{0}]_{j}$ is either the square of a standard Gaussian random variable, $\sqrt{2}$ times the product of two independent standard Gaussian random variables, or one. Hence, $f_{0}]_{j}$ is subexponential with mean and $\|[f_{0}]_{j}\|_{\psi_{1}}$ both bounded by $\mathcal{O}(1)$. As a result, the product $f_{0}]_{j}e$ is $\frac{1}{2}$-sub-Weibull, with the sub-Weibull norm being $\mathcal{O}(E)$. By (Hao et al., 2019, Theorem 3.1),
@@ 4.3 Matrix factorization bound
Given two $m\times n$ matrices $A,B$, we are interested in bounding $\min_{S^{\top}S=I}\|SA-B\|_{F}$ using $\|A^{\top}A-B^{\top}B\|_{F}$. The minimum problem is known as the orthogonal Procrustes problem, solved in (Schoenemann, 1964; Schönemann, 1966). Specifically, the minimum is attained at $S=UV^{\top}$, where $U\Sigma V^{\top}=BA^{\top}$ is its singular value decomposition.
If $m\leq n$ and $\mathrm{rank}(A)=m$, then the following lemma from (Tu et al., 2016) establishes that the distance between $A$ and $B$ is of the same order of $\|A^{\top}A-B^{\top}B\|_{F}$.
Lemma 3 ((Tu et al., 2016, Lemma 5.4)).
@@ 4.4 Perturbed linear regression bound
By (Wainwright, 2019, Theorem 6.1), the Gaussian ensemble $G$ satisfies that with probability at least 1p1-p,
@@
where we consider $\epsilon_{x}$ and $\epsilon_{y}$ as quantities much smaller than one such that terms like $\epsilon_{x}^{2},\epsilon_{x}\epsilon_{y}$ are absorbed into $\epsilon_{x},\epsilon_{y}$. It remains to control $\|G^{\dagger}E\|_{2}$. It is proved in (Mhammedi et al., 2020, Section B.2.11) via a covering number argument that with probability at least 1p1-p,
@@ 4.4.1 Proof of Claim 1
Similar to the proof of Lemma 5, by (Mhammedi et al., 2020, Section B.2.11), with probability at least 1p1-p,
@@ 4.5.3 Proof of Claim 4
The following lemma is the counterpart of (Fazel et al., 2018, Lemma 10) in the LTV setting with system noises.
@@
This problem is studied in (Mania et al., 2019) for the infinite-horizon LTI setting; here we extend their result to the finite-horizon LTV setting.
We adopt the following compact formulation of a finite-horizon LTV system, as introduced in (Zhang et al., 2021, Section 3), to reduce our setting to the infinite-horizon LTI one:
@@
Let $\hat{P}=\textnormal{diag}(\hat{P}_{0},\ldots,\hat{P}_{T})$ be the optimal cumulative cost matrices in system $(\hat{A},\hat{B},\hat{Q},R^{\ast})$ by the RDE (2.6). With a slight abuse of notation, define $K^{\ast}:=[\textnormal{diag}(K^{\ast}_{0},\ldots,K^{\ast}_{T-1}),0_{d_{u}T\times d_{x}}]$, where $(K^{\ast}_{t})_{t=0}^{T-1}$ is the optimal controller in system $(A^{\ast},B^{\ast},Q^{\ast},R^{\ast})$, and define $\hat{K}$ similarly for system $(\hat{A},\hat{B},\hat{Q},R^{\ast})$. By definition, $A^{\ast}$ is stable (hence, stabilizable), $(A^{\ast},Q^{\ast})$ is observable in the sense of LTI systems, and $R^{\ast}\succ 0$. Therefore, by (Mania et al., 2019, Propositions 1 and 2), there exists a dimension-free constant $\varepsilon_{0}>0$ with $\varepsilon_{0}^{-1}$ depending polynomially on problem parameters such that as long as $\varepsilon\leq\varepsilon_{0}$,
@@
Finally, we are ready to present the main lemma of this section below, which combines Lemmas 7 to 9 to provide an end-to-end guarantee for the certainty equivalent optimal controller in linear quadratic control with both rank-deficient and full-rank stages.
@@
For any feedback gains $(K_{t})_{t=0}^{T-1}$, let $J(K)$ denote the expected cumulative cost in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1}$, $(Q^{\ast}_{t})_{t=0}^{T})$ under state feedback control $u_{t}=K_{t}x_{t}$ for $t\geq 0$, and $J^{\delta}(K)$ denote the expected cumulative cost in system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$ under control $u_{t}=K_{t}(x_{t}+\delta_{t})$, where $\delta_{t}$ is a perturbation, for $t\geq 0$. Under control $u_{t}=K_{t}(x_{t}+\delta_{t})$ for all $t\geq 0$, let $\Xi_{t}(K):=\mathrm{Cov}(x_{t})$ in the system $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(Q^{\ast}_{t})_{t=0}^{T})$, and assume that $\delta_{t}$ follows an arbitrary zero-mean Gaussian distribution, can be correlated with $x_{t}$, and satisfies $\|\mathrm{Cov}(\delta_{t})\|_{2}^{1/2}\leq\varphi_{1}\max_{\tau\leq t}\|\Xi_{\tau}(K)\|_{2}^{1/2}$ for $0\leq t\leq\ell-1$ and $\|\mathrm{Cov}(\delta_{t})\|_{2}^{1/2}\leq\varphi_{2}\max_{\tau\leq t}\|\Xi_{\tau}(K)\|_{2}^{1/2}$ for $t\geq\ell$.
There exist dimension-free constants $\kappa,a_{2},a_{3},a_{4}>1$ that depend polynomially on $c$ and other problem parameters and dimension-free constant $\varepsilon_{0}>0$ with $\varepsilon_{0}^{-1}$ depending polynomially on problem parameters, such that under the conditions that 1) $\varepsilon_{1},\varphi_{1}$ are small enough to ensure $\|\hat{B}_{t}\|_{2},\|\hat{Q}_{t}\|_{2}$ are of order $\mathcal{O}(1)$, $\lambda_{\min}(\hat{Q}_{t})=\Omega(1)$ for $0\leq t\leq\ell-1$ and $\varepsilon_{1}+\varphi_{1}=\mathcal{O}((\kappa a_{2}+a_{3}^{2})^{-T})$, and that 2) $\varepsilon_{2},\varphi_{2}$ are small enough to ensure $\varepsilon_{2}\leq\varepsilon_{0}$ and $\varepsilon_{2}+\varphi_{2}=\mathcal{O}(\varepsilon_{1}+\varphi_{1})=\mathcal{O}({T}^{-1})$, we have
@@
$J_{2}(\hat{K})=\mathbb{E}\big[\|x_{\ell}^{\delta}\|_{P_{\ell}^{\hat{K}}}^{2}\big]+\sum\nolimits_{t=\ell}^{T-1}\mathbb{E}[w_{t}^{\top}P_{t+1}^{\hat{K}}w_{t}].$